# Make LoRA Great Again: Boosting LoRA with Adaptive Singular Values and Mixture-of-Experts Optimization Alignment

### v2

@@ Abstract
While Low-Rank Adaptation (LoRA) enables parameter-efficient fine-tuning for Large Language Models (LLMs), its performance often falls short of Full Fine-Tuning (Full FT). Current methods optimize LoRA by initializing with static singular value decomposition (SVD) subsets, leading to suboptimal leveraging of pre-trained knowledge. Another path for improving LoRA is incorporating a Mixture-of-Experts (MoE) architecture. However, weight misalignment and complex gradient dynamics make it challenging to adopt SVD prior to the LoRA MoE architecture. To mitigate these issues, we propose Great LoRA Mixture-of-Expert (GOAT), a framework that (1) adaptively integrates relevant priors using an SVD-structured MoE, and (2) aligns optimization with full fine-tuned MoE by deriving a theoretical scaling factor. We demonstrate that proper scaling, without modifying the architecture or training algorithms, boosts LoRA MoE’s efficiency and performance. Experiments across 25 datasets, including natural language understanding, commonsense reasoning, image classification, and natural language generation, demonstrate GOAT’s state-of-the-art performance, closing the gap with Full FT.
@@ 1 Introduction
Recent large language models (LLMs) have shown impressive capabilities (Dai et al., 2024; Touvron et al., 2023; Yang et al., 2024; OpenAI et al., 2024), but fine-tuning them for downstream tasks is computationally expensive (Hu et al., 2021; Zhao et al., 2024). To reduce costs, parameter-efficient fine-tuning (PEFT) techniques (Hu et al., 2021; Pfeiffer et al., 2021; Houlsby et al., 2019; Tian et al., 2024) have been proposed. Among them, LoRA (Hu et al., 2021) is popular for its simplicity and effectiveness. It reparameterizes the weight matrix $W\in{\mathbb{R}}^{m\times n}$ into $W=W_{0}+BA$, where $W_{0}\in{\mathbb{R}}^{m\times n}$ is a frozen full-rank matrix, and $B\in\mathbb{R}^{m\times r}$,$A\in\mathbb{R}^{r\times n}$ are low-rank adapters to be learned. Since the rank $r\ll\min(m,n)$, LoRA only updates a small fraction of the parameters, greatly reducing memory usage.
Despite its computational efficiency, LoRA often underperforms full fine-tuning (Full FT) (Wang et al., 2024d, c; Fan et al., 2024), even with Mixture-of-Experts (MoE) architectures  (Zadouri et al., 2024; Liu & Luo, 2024; Tian et al., 2024). Our rigorous analysis identifies two key factors limiting LoRA’s performance: (1) Suboptimal Initialization: The isotropic random initialization for matrix $A$ and zero initialization for matrix $B$ provide a non-informative prior, resulting in unguided optimization subspaces. While Wang et al. (2024b); Meng et al. (2024) applied singular value decomposition (SVD) for better initialization, their reliance on a static, predefined subset of pre-trained weights limits the capture of the full range of pre-trained knowledge. It raises the question: Can we adaptively integrate relevant priors of pre-trained knowledge based on input? (2) Unaligned Optimization: Furthermore, the intrinsic low-rank property of LoRA leads to large gradient gaps and slow convergence in optimization, therefore underperforming Full FT. In LoRA MoE scenarios, the total rank is split among experts, resulting in lower ranks and further increasing this challenge. Existing strategies (Wang et al., 2024c, d) focus only on single LoRA architectures and ignore the added complexity of random top-$k$ routing and multiple expert weights within MoE architecture. When SVD-based initialization is applied to LoRA MoE, weight alignment becomes a challenge, which has never been considered in previous methods that used zero initialization. This further raises the question: How do we mitigate the optimization gap in LoRA MoE initialized with prior information?
@@ 2.1 Rethinking Singular-Value Initialization
We observe PiSSA freezes minor singular values and fine-tunes only the components $U_{0}\Sigma_{0}V_{0}^{\top}$ with the largest norms, achieving the optimal approximation to $W_{0}$. [^3] Proof in Appendix A In contrast, MiLoRA and KaSA retain segment $0\sim(l-1)$ as preserved pretrained knowledge, but KaSA treats the minor $U_{l}\Sigma_{l}V_{l}^{\top}$ as noise and replaces it with a new random $U^{\text{r}}\Sigma^{\text{r}}V^{\text{r}\top}$. In practice, PiSSA converges faster by focusing on principal singular values, while MiLoRA and KaSA preserve more pre-trained knowledge for better final performance.
@@ 3.1 LoRA MoE Architecture
We integrate LoRA into the MoE framework, retaining the router (Equation (10)) and using the balance loss from vanilla MoE [^4] See Appendix B. Each expert $W^{i}$ is replaced by low-rank matrices $B^{i}\in\mathbb{R}^{m\times d}$ and $A^{i}\in\mathbb{R}^{d\times n}$, where $d=\frac{r}{E}$:
@@ 3.2 Adaptive Priors Initialization
where $t=\frac{\min(m,n)}{E},k=(j-1)t$ is the starting index from segement for $j$-th expert, $d=\frac{r}{E}$ is each expert rank. Then we can construct each expert by $\left(U^{\prime},\Sigma^{\prime},V^{\prime\top}\right)\in\mathcal{E}_{r}$:
@@ 3.3 Theoretical Optimization Alignment
Directly applying SVD priors in MoE architectures causes weight misalignment and complex gradient dynamics, a challenge not encountered with previous zero initialization methods. Moreover, the gap in MoE-based architectures remains under-explored. We derive the following theorems to address this and show how scaling resolves the issue.
@@
Theorem 3.1 addresses the performance gap in single LoRA architectures (Wang et al., 2024c, d). For MoE architectures, however, routers and top-$k$ selection complicate direct alignment. Thus, we focus on Full FT MoE and establish:
@@
As $n\gg r$, it is typically the case that $s>2$, which explains why standard scaling is insufficient and why simple scaling enhances effectiveness, as demonstrated in Section 2.2. While it is tricky to directly analyze complex gradient dynamics with SVD priors, an alternative approach is recognizing that larger scaling $s$ and $\rho$ ensure $B$ and $A$ become small and approach zero (Equation (19)), aligning with the settings in Theorem 3.5. Thus, we adopt this scaling factor in GOAT, and in practice, this approximation performs well (see Section 4.3). For scenarios with proper scaling, we extend the method to “GOAT+”, as detailed in Appendix D.
@@ 4.1 Baselines
For a fair comparison, we closely follow the configurations from prior studies (Hu et al., 2021; Meng et al., 2024; Wang et al., 2024d). Details on the baselines are in Appendix E.
@@ 4.2 Datasets
    Commonsense Reasoning (CR): We fine-tune LLaMA2-7B on Commonsense170K and evaluate on 8 commonsense reasoning datasets  (Hu et al., 2023).
@@
Due to the 8x memory requirements of Full FT MoE, we only evaluate it on IC and NLU tasks. Detailed of the datasets can be found in Appendix E.1.
@@ 4.4 Ablation Study
We conduct ablation experiments to evaluate the impact of our adaptive priors initialization and gradient scaling, as summarized in Table 5. Our initialization, with or without MoE scaling, consistently outperforms other methods [^5] Details provided in Appendix E.3 (note that no SVD-based initialization corresponds to the original zero initialization, yielding 81.06/80.26). Without MoE, initializing a single LoRA with our SVD fragments achieves a performance of 77.62. In contrast, our MoE architecture achieves 80.35, demonstrating its clear advantage in effectively integrating expert functionalities.
@@ 4.9 Computation Analysis
The “# Params (%)” column in Tables 1, 2, 3, and 4 compares the parameter ratios of LoRA baselines and GOAT to full fine-tuning MoE. GOAT achieves state-of-the-art performance with a parameter size of $O(Hr)+O(He)$, significantly smaller than Full FT’s $O(H^{2})$ and Full FT MoE’s $O(kH^{2})$. Since $r,e\ll H$, GOAT is much more efficient. Detailed analysis is in Appendix F.1.
@@
To compare with Full FT MoE, we estimate the memory usage, runtime, and performance of FT MoE based on the single GPU runtime of Full FT. As shown in Table 7, the LoRA-MoE series trains much faster than Full FT MoE. Among LoRA-MoE variants, our method achieves the best performance with identical memory and time costs. FLOPs analysis (see Appendix F.2) reveals that Full FT MoE scales as $O(ksH^{2})$, while LoRA MoE simplifies to $O(sH^{2})$ since $k<e$ and $r\ll H$. Thus, LoRA MoE’s FLOPs remain nearly constant, independent of $k$, unlike Full FT MoE, which scales linearly with $k$.
@@ 6 Conclusion
7 Impact Statements
@@
Appendix A Proof related with PiSSA Select Segment
Lemma A.1.
@@ Appendix A Proof related with PiSSA Select Segment
Appendix B Load Balance Loss
@@
Appendix C Proof of Theoretical Results
C.1 Proof of Lemma 2.2
@@ Appendix C Proof of Theoretical Results
Given that $\mathbb{E}_{A_{0}}[A_{0}^{\top}A_{0}]=\frac{r}{3n}{\mathbf{I}}_{n\times n}$ (use Leaky ReLU with negative slope $\sqrt{5}$, that is $\text{Var}(A)=\frac{1}{3n}$), we can get $s=\sqrt{\frac{3n\eta}{r}}$
@@
Appendix D Extend Our Method to Scenarios with Proper Scaling
@@
$s_{i}=s_{i}\frac{\sqrt{\sigma_{0}}}{\sqrt{\sigma_{i}}}$ (83)
@@
We refer to this extended method as GOAT+, and its performance across all benchmarks is presented in Table 8. While designed for different scenarios, it demonstrates performance comparable to GOAT.
@@
GOAT+ 30.54 89.61 81.54 82.41 71.02
Appendix E Experiment Details
E.1 Dataset details

### v3

@@ Abstract
While Low-Rank Adaptation (LoRA) enables parameter-efficient fine-tuning for Large Language Models (LLMs), its performance often falls short of Full Fine-Tuning (Full FT). Current methods optimize LoRA by initializing with static singular value decomposition (SVD) subsets, leading to suboptimal leveraging of pre-trained knowledge. Another path for improving LoRA is incorporating a Mixture-of-Experts (MoE) architecture. However, weight misalignment and complex gradient dynamics make it challenging to adopt SVD prior to the LoRA MoE architecture. To mitigate these issues, we propose Great LoRA Mixture-of-Expert (GOAT), a framework that (1) adaptively integrates relevant priors using an SVD-structured MoE, and (2) aligns optimization with full fine-tuned MoE by deriving a theoretical scaling factor. We demonstrate that proper scaling, without modifying the architecture or training algorithms, boosts LoRA MoE’s efficiency and performance. Experiments across 25 datasets, including natural language understanding, commonsense reasoning, image classification, and natural language generation, demonstrate GOAT’s state-of-the-art performance, closing the gap with Full FT. Our code is available at: https://github.com/Facico/GOAT-PEFT.
@@ 1 Introduction
Recent large language models (LLMs) have shown impressive capabilities (Dai et al., 2024; Touvron et al., 2023; Yang et al., 2024; OpenAI et al., 2024), but fine-tuning them for downstream tasks is computationally expensive (Hu et al., 2021; Zhao et al., 2024). To reduce costs, parameter-efficient fine-tuning (PEFT) techniques (Hu et al., 2021; Pfeiffer et al., 2021; Houlsby et al., 2019; Tian et al., 2024; Fan et al., 2024b) have been proposed. Among them, LoRA (Hu et al., 2021) is popular for its simplicity and effectiveness. It reparameterizes the weight matrix $W\in{\mathbb{R}}^{m\times n}$ into $W=W_{0}+BA$, where $W_{0}\in{\mathbb{R}}^{m\times n}$ is a frozen full-rank matrix, and $B\in\mathbb{R}^{m\times r}$,$A\in\mathbb{R}^{r\times n}$ are low-rank adapters to be learned. Since the rank $r\ll\min(m,n)$, LoRA only updates a small fraction of the parameters, greatly reducing memory usage.
Despite its computational efficiency, LoRA often underperforms full fine-tuning (Full FT) (Wang et al., 2024d, c; Fan et al., 2024a), even with Mixture-of-Experts (MoE) architectures  (Zadouri et al., 2024; Liu & Luo, 2024; Tian et al., 2024). Our rigorous analysis identifies two key factors limiting LoRA’s performance: (1) Suboptimal Initialization: The isotropic random initialization for matrix $A$ and zero initialization for matrix $B$ provide a non-informative prior, resulting in unguided optimization subspaces. While Wang et al. (2024b); Meng et al. (2024) applied singular value decomposition (SVD) for better initialization, their reliance on a static, predefined subset of pre-trained weights limits the capture of the full range of pre-trained knowledge. It raises the question: Can we adaptively integrate relevant priors of pre-trained knowledge based on input? (2) Unaligned Optimization: Furthermore, the intrinsic low-rank property of LoRA leads to large gradient gaps and slow convergence in optimization, therefore underperforming Full FT. In LoRA MoE scenarios, the total rank is split among experts, resulting in lower ranks and further increasing this challenge. Existing strategies (Wang et al., 2024c, d) focus only on single LoRA architectures and ignore the added complexity of random top-$k$ routing and multiple expert weights within MoE architecture. When SVD-based initialization is applied to LoRA MoE, weight alignment becomes a challenge, which has never been considered in previous methods that used zero initialization. This further raises the question: How do we mitigate the optimization gap in LoRA MoE initialized with prior information?
@@ 2.1 Rethinking Singular-Value Initialization
We observe PiSSA freezes minor singular values and fine-tunes only the components $U_{0}\Sigma_{0}V_{0}^{\top}$ with the largest norms, achieving the optimal approximation to $W_{0}$. [^3] Proof in Appendix B In contrast, MiLoRA and KaSA retain segment $0\sim(l-1)$ as preserved pretrained knowledge, but KaSA treats the minor $U_{l}\Sigma_{l}V_{l}^{\top}$ as noise and replaces it with a new random $U^{\text{r}}\Sigma^{\text{r}}V^{\text{r}\top}$. In practice, PiSSA converges faster by focusing on principal singular values, while MiLoRA and KaSA preserve more pre-trained knowledge for better final performance.
@@ 3.1 LoRA MoE Architecture
We integrate LoRA into the MoE framework, retaining the router (Equation (10)) and using the balance loss from vanilla MoE [^4] See Appendix C. Each expert $W^{i}$ is replaced by low-rank matrices $B^{i}\in\mathbb{R}^{m\times d}$ and $A^{i}\in\mathbb{R}^{d\times n}$, where $d=\frac{r}{E}$:
@@ 3.2 Adaptive Priors Initialization
where $t=\frac{\min(m,n)}{E},k=(j-1)t$ is the starting index from segment for $j$-th expert, $d=\frac{r}{E}$ is each expert rank. Then we can construct each expert by $\left(U^{\prime},\Sigma^{\prime},V^{\prime\top}\right)\in\mathcal{E}_{r}$:
@@ 3.3 Theoretical Optimization Alignment
Directly applying SVD priors in MoE architectures causes weight misalignment and complex gradient dynamics, a challenge not encountered with previous zero initialization methods. Moreover, the gap in MoE-based architectures remains under-explored. We derive the following theorems to mitigate this and show how scaling resolves the issue.
@@
Theorem 3.1 mitigates the performance gap in single LoRA architectures (Wang et al., 2024c, d). For MoE architectures, however, routers and top-$k$ selection complicate direct alignment. Thus, we focus on Full FT MoE and establish:
@@
As $n\gg r$, it is typically the case that $s>2$, which explains why standard scaling is insufficient and why simple scaling enhances effectiveness, as demonstrated in Section 2.2. While it is tricky to directly analyze complex gradient dynamics with SVD priors, an alternative approach is recognizing that larger scaling $s$ and $\rho$ ensure $B$ and $A$ become small and approach zero (Equation (19)), aligning with the settings in Theorem 3.5. Thus, we adopt this scaling factor in GOAT, and in practice, this approximation performs well (see Section 4.3). For scenarios with proper scaling, we extend the method to “GOAT-s”, as detailed in Appendix E.
@@ 4.1 Baselines
For a fair comparison, we closely follow the configurations from prior studies (Hu et al., 2021; Meng et al., 2024; Wang et al., 2024d). Details on the baselines are in Appendix G.
@@ 4.2 Datasets
    Commonsense Reasoning (CR): We fine-tune LLaMA2-7B on Commonsense170K and evaluate on 8 commonsense reasoning datasets  (Hu et al., 2023) (multi-domain setting).
@@
Due to the huge memory requirements of Full FT MoE, we only evaluate it on IC and NLU tasks. Detailed of the datasets can be found in Appendix G.1.
@@ 4.4 Ablation Study
We conduct ablation experiments to evaluate the impact of our adaptive priors initialization and gradient scaling, as summarized in Table 5. Our initialization, with or without MoE scaling, consistently outperforms other methods [^5] Details provided in Appendix G.3 (note that no SVD-based initialization corresponds to the original zero initialization, yielding 81.06/80.26). Without MoE, initializing a single LoRA with our SVD fragments achieves a performance of 77.62. In contrast, our MoE architecture achieves 80.35, demonstrating its clear advantage in effectively integrating expert functionalities.
@@ 4.9 Computation Analysis
The “# Params (%)” column in Tables 1, 2, 3, and 4 compares the parameter ratios of LoRA baselines and GOAT to full fine-tuning MoE. GOAT achieves state-of-the-art performance with a parameter size of $O(Hr)+O(He)$, significantly smaller than Full FT’s $O(H^{2})$ and Full FT MoE’s $O(kH^{2})$. Since $r,e\ll H$, GOAT is much more efficient. Detailed analysis is in Appendix H.1.
@@
To compare with Full FT MoE, we estimate the memory usage, runtime, and performance of FT MoE based on the single GPU runtime of Full FT. As shown in Table 7, the LoRA-MoE series trains much faster than Full FT MoE. Among LoRA-MoE variants, our method achieves the best performance with identical memory and time costs. FLOPs analysis (see Appendix H.2) reveals that Full FT MoE scales as $O(ksH^{2})$, while LoRA MoE simplifies to $O(sH^{2})$ since $k<e$ and $r\ll H$. Thus, LoRA MoE’s FLOPs remain nearly constant, independent of $k$, unlike Full FT MoE, which scales linearly with $k$.
@@ 6 Conclusion
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant No. 62276110, No. 62172039 and in part by the fund of Joint Laboratory of HUST and Pingan Property & Casualty Research (HPL). We thank the Shanghai AI Laboratory for supporting GPU resources. The authors would also like to thank the anonymous reviewers for their comments on improving the quality of this paper.
Impact Statement
@@ Acknowledgments
Appendix A Pseudocode
Algorithm 1 GOAT
0:  Input vector $x$, input dimension $n$, hyperparameters $\eta,\rho$, number of experts $E$
0:  Output $y=\tilde{W}_{0}(x)+\sum_{i=1}^{E}w^{i}(x)\,s\,B_{0}^{i}A_{0}^{i}(x)$
1:  function Initialization
2:        Scaling factor: $s\leftarrow\sqrt{3n\eta/\rho}$
3:        SVD decomposition: $W_{0}=U\,\Sigma\,V^{\top}$
4:         for $i=1$ to $E$ do
5:            $B_{0}^{i}\leftarrow\sqrt{1/(s\rho)}\,U^{\prime}\Sigma^{\prime 1/2}$
6:            $A_{0}^{i}\leftarrow\sqrt{1/(s\rho)}\,\Sigma^{\prime 1/2}V^{\prime\top}$
7:         end for
8:        $W_{\text{res}}^{+}\leftarrow\frac{s}{E}\sum_{i=1}^{E}B_{0}^{i}A_{0}^{i}$
9:        $\tilde{W}_{0}\leftarrow W_{0}-W_{\text{res}}^{+}$
10:        return: $\tilde{W}_{0},\{B_{0}^{i},A_{0}^{i}\}$
11:  end function
1:  function Forward($x$
2:        Compute gating weight $w^{i}(x)$ ($1\leq i\leq E$)
3:        return: $\tilde{W}_{0}(x)+\sum_{i=1}^{E}w^{i}(x)\,s\,B_{0}^{i}A_{0}^{i}(x)$
4:  end function
Appendix B Proof related with PiSSA Select Segment
Lemma B.1.
@@ Appendix B Proof related with PiSSA Select Segment
Appendix C Load Balance Loss
@@
Appendix D Proof of Theoretical Results
D.1 Proof of Lemma 2.2
@@ Appendix D Proof of Theoretical Results
Given that $\mathbb{E}_{A_{0}}[A_{0}^{\top}A_{0}]=\frac{r}{3n}{\mathbf{I}}_{n\times n}$ (use Leaky ReLU (Xu et al., 2015) with negative slope $\sqrt{5}$, that is $\text{Var}(A)=\frac{1}{3n}$), we can get $s=\sqrt{\frac{3n\eta}{r}}$
@@
Appendix E Extend Our Method to Scenarios with Proper Scaling
@@
$s_{i}=s_{1}\frac{\sqrt{\sigma_{0}}}{\sqrt{\sigma_{i}}}$ (83)
@@
We refer to this extended method as GOAT-s, and its performance across all benchmarks is presented in Table 8. While designed for different scenarios, it demonstrates performance comparable to GOAT.
@@
GOAT-s 30.54 89.61 81.54 82.41 71.02
Appendix F Further Discussion
F.1 Discussion on the Practical Applications and Real-world Impact of LoRA MoE
MoE is popular for managing large parameters while activating only a sparse subset during inference, making it ideal for large-scale models. However, in Section 4.9 and Table 7, without optimization, fully fine-tuning an MoE model significantly increases trainable parameters and FLOPs compared to Full FT.
LoRA MoE addresses these challenges by replacing experts with low-rank matrices, reducing computation, preserving MoE benefits, and enabling faster training, lower memory usage, and reduced energy consumption—crucial for resource-limited or real-time applications.
For instance, in NLP, where large-scale models are common, LoRA MoE achieves SOTA performance with lower computational cost. This efficiency benefits industries like autonomous driving, healthcare (Tian et al., 2024), where lower latency and costs enhance performance and scalability.
Overall, LoRA MoE balances MoE’s model capacity with cost-effective deployment, making it adaptable to various real-world applications.
F.2 Analysis of the Phenomenon Where PiSSA and MiLoRA May Perform Worse Than LoRA in Experiments
Previous works (Wang et al., 2024b, a) show that PiSSA and MiLoRA don’t always outperform LoRA. KaSA found that PiSSA accelerates convergence but uses limited pre-trained knowledge at lower ranks, limiting performance. Similarly, MiLoRA’s minimal adjustments to pre-trained weights often fail to improve over LoRA. In Table 1, we adopt the same rank settings as KaSA and reach the same conclusion.
In contrast, our method consistently achieves superior performance across both low and high ranks by effectively balancing convergence speed and final performance, as demonstrated in Tables 1,2,3,4 and Figure 5.
F.3 Compared to Other Routing Techniques
As shown in Table 9, we extend our experiments to include alternative routing strategies such as top-p routing and a top‑k variant with shared experts. We find that, compared to other approaches, setting $k=2$ achieves the best performance. We will incorporate these into our revised version of the paper.
Table 9: Comparison of Routing Strategies on Average Accuracy
Routing Strategy Avg. ACC (%)
Ours (top-$k=2$) 81.49
Top-$p$ ($p=0.25$) 79.40
Top-$k$ + Shared Expert 78.68
F.4 Analysis of the coefficient for the balance loss
We use top-k routing with k=2 and set the coefficient for the balance loss to 1e-3. As shown in Table 10, we attach the load-balancing loss coefficient experiment by activating 2 out of 8 experts on Cars.
Table 10: Performance Comparison under Different Coefficient Values
Coefficient GOAT MoLoRA HydraLoRA
$1\times 10^{-1}$ 49.09 49.02 48.45
$1\times 10^{-2}$ 50.52 49.33 49.45
$1\times 10^{-3}$ 53.50 50.83 48.42
$1\times 10^{-4}$ 51.53 49.03 48.52
00 49.85 48.02 49.06
We can observe that setting the coefficient too low (e.g., 0 or 1e-4) leads to expert imbalances, which in turn degrades performance. Conversely, excessively high coefficients (e.g., 0.01 or 0.1) can disrupt the normal learning process. Our results show that a coefficient of 1e-3 achieves the best tradeoff in GOAT/MoLoRA between balancing expert load and maintaining stable learning.
Notably, GOAT consistently outperforms across all tested coefficients, demonstrating its robustness in these settings.
F.5 Is expert activation balanced without the coefficient for the balance loss?
Table 11: Expert Activation Distribution of GOAT without Load Balancing
Method f1 f2 f3 f4 f5 f6 f7 f8
GOAT w/o Load Balance 0.1043 0.1379 0.1275 0.1094 0.1207 0.1405 0.1259 0.1338
We further ablate the load-balancing component in GOAT (2in8, Cars task) and observe that all experts remain active (see Table 11), indicating that each SVD chunk contributes meaningfully to the final representation.
Appendix G Experiment Details
G.1 Dataset details